# A Computational Tsirelson’s Theorem for the Value of Compiled XOR Games

### v2

@@ 1 Introduction
In a nonlocal game, a polynomially-bounded classical verifier interacts with multiple non-communicating provers. A set of predetermined rules governs whether the provers win or lose the game. The classical value of a nonlocal game is the highest winning probability that any classical (or unentangled) provers can achieve, whereas the quantum value is the highest winning probability that any quantum (potentially entangled) provers can achieve. Although the study of nonlocal games originated with Bell’s theorem [Bel64], they have become central in the study of quantum complexity theory through their connection with multiprover interactive proofs [CHTW04]. This connection has led to many protocols for efficiently certifying classical and quantum computations in the two-prover setting [RUV13, CGJV20, Gri17, JNV+21].
A central question in the nonlocal game paradigm is whether the techniques and results can be translated to a model where a verifier interacts with a single prover. The single-prover setting is theoretically appealing and practically motivated since a non-communication assumption between provers can be challenging to enforce. Recently, Kalai et al. [KLVY23] proposed a generic procedure that compiles any $k$-player one-round nonlocal game into a multi-round single-prover protocol, using a quantum homomorphic encryption scheme [Mah20, Bra18]. In the same work, the authors showed that the compilation process preserves the classical value of the nonlocal game. Shortly after, Natarajan and Zhang [NZ23] proved that the same compiler preserves the quantum value of the $2$-player CHSH game [CHSH69]. This was sufficient to achieve efficient delegation of BQP computations, however, their work left open the question of whether the Kalai et al. compiler preserved the quantum value for a more general class of games.
A natural generalization of the CHSH game is the class of two-player XOR games [CHTW04], where the winning predicate depends on the XOR of the players’ single-bit answers. These games have a foundational place in the theory of nonlocal games thanks to an influential theorem of Tsirelson [Tsi80, Tsi87], which exactly characterizes their optimal quantum strategies in terms of a semidefinite program. This makes XOR games one of the few classes of nonlocal games where the optimal quantum strategies are well understood, and finding the quantum value is computationally tractable. As such, they are a convenient stepping stone towards understanding more general families of nonlocal games.
@@ 1.1 Results
Our main result is that the quantum value of any two-player XOR game is preserved by the compilation procedure of [KLVY23], resolving a question posed in [NZ23]. We give an informal statement of our result; a more detailed statement can be found in Theorem 3.1.
@@
The negligible factors in Theorem 1.1 come from the fact that the compiled game assumes the existence of quantum homomorphic encryption, whose security is computational and depends on chosen parameters. It is well-known that quantum homomorphic encryption can be constructed from the hardness of the standard learning with errors (LWE) problem [Mah18, Bra18], and thus our theorem holds up to the same assumption. A useful consequence of this result is that we can apply the extensive literature on XOR games to give new examples of well-behaved compiled games. For instance, we obtain the following new results as almost immediate corollaries of the proof of our main theorem:
  • Self-testing for all XOR games Using a result from Slofstra [Slo11], we show that every compiled XOR game self-tests some properties of the provers’ measurement operators: in particular, for each XOR game there is a set of relations between the Alice and Bob operators that are approximately satisfied for any near-optimal strategy for the compiled game (Theorem 3.6). As a special case, we recover the self-test for anticommutation in [NZ23].
  • Parallel repetition Building on [CSUU08], we show that the quantum value of compiled parallel repeated XOR games satisfies perfect exponential decay in the number of parallel repetitions (Theorem 3.11). This is the first parallel repetition theorem for compiled games.
@@
  • Self-testing for the Magic Square game: We show that the compilation of the Magic Square game of Mermin and Peres [Mer90, Per90, Ara02] is a self-test for a pair of anticommuting operators (Theorem 5.7). This mirrors the statement in the two-prover settings [WBMS16].
@@ 1.2 Technical Outline
Non-local games and the [KLVY23] compiler.
@@
It is easy to show that the compiled value is always at least the nonlocal value, as any nonlocal strategy can be implemented using the homomorphic encryption scheme. It is far less obvious how to establish upper bounds on the compiled value. This is because all known general techniques for upper-bounding the value of a nonlocal game make essential use of the spatial separation between the two provers. Nevertheless, we are able to show that the two specific techniques used by [NZ23] to analyze the compiled CHSH game both generalize to all compiled XOR games. We believe the two approaches are complementary and improve our understanding of compiled nonlocal games from dual perspectives.
The macroscopic locality approach.
Our first approach generalizes the “macroscopic locality” proof of soundness of [NZ23] for the CHSH game, which is based on arguments contained in [Roh14]. One can use the IND-CPA security of the homomorphic encryption to upper-bound the maximum possible probability with which any efficiently measurable Bob operator can guess the value of Alice’s question $x$. Intuitively, this allows us to establish that Bob’s strategy cannot depend, in any detectable way, on Alice’s question. Applying Jensen’s inequality to Bob’s operator, we can then bound, up to a negligible function, the bias of every quantum strategy for the compiled XOR game by the bias of some vector strategy for the nonlocal XOR game. These vector strategies can be thought of as a vector solution to the semidefinite program which computes the bias of the XOR game. By Tsirelson’s theorem, every such vector strategy gives rise to a quantum strategy for the nonlocal XOR game, yielding the desired bound on the winning probability for the compiled game in terms of the optimal nonlocal value of the original XOR game. As a corollary, we obtain a one-sided rigidity or “self-testing” statement for any compiled XOR game: For any near-optimal strategy, it must hold that $A_{x}|\psi\rangle\approx\sum_{y}c_{xy}B_{y}A_{x}|\psi\rangle$ for explicit coefficients $c_{xy}$.
The sum of squares approach.
@@
A more abstract approach to upper-bounding the quantum value of a nonlocal game is to associate with a game a formal polynomial $p$, whose variables are identified with the observables $A_{x}$ and $B_{y}$ so that $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ is precisely the winning probability of the strategy. (For example, in the case of CHSH, the polynomial is $p(A_{x},B_{y})=\frac{1}{2}+\frac{1}{8}(A_{0}B_{0}+A_{0}B_{1}+A_{1}B_{0}-A_{1}B_{1})$, and it can be easily verified that $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ captures the winning probability of the strategy ($(\{A_{x}\},\{B_{y}\},|\psi\rangle)$ for the CHSH game. Moreover, to prove that the quantum value of a game is at most $\omega$, it suffices to give a sum-of-squares (SOS) certificate demonstrating that $\omega-p(A_{x},B_{y})$ is a positive operator whenever $p$ is evaluated on variables that corresponds to valid observables. Such a certificate consists of a decomposition of $\omega-p(A_{x},B_{y})$ as a positive operator
$\omega-p(A_{x},B_{y})=\sum_{i}q_{i}(A_{x},B_{y})^{*}q_{i}(A_{x},B_{y})+g(A_{x},B_{y}),$
@@
where the $q_{i}$’s and $g$ are polynomials, and $g$ is a polynomial that evaluates to zero whenever the variables are valid observables. In particular, $g$ lies in the ideal generated by a certain collection of polynomial constraints on the variables $A_{x}$ and $B_{y}$, including the commutator constraint $[A_{x},B_{y}]=0$ for all $x,y$, which formally represents the spatial separation of the players.
@@
The reason why such an SOS certificate suffices to upper bound the game value in the nonlocal case is that (1) taking the expectation $\langle\psi|\omega-p(A_{x},B_{y})|\psi\rangle$ maps positive operators to positive numbers, and (2) $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ is equal to the winning probability of strategy $S=(\{A_{x},B_{y}\},|\psi\rangle)$. Property (1) gives that evaluating the expectation on $\omega-p(A_{x},B_{y})$ yields a positive number, and property (2) says that this positive number is equal to $\omega-p_{win}$, where $p_{win}$ is the winning probability of the strategy $S=(\{A_{x},B_{y}\},|\psi\rangle)$. Therefore, in the nonlocal case, taking the expectation plays the role of what is often known as a “pseudo-expectation” in the SOS literature.
@@
In general, an SOS certificate of the aforementioned form does not automatically yield an upper bound on the compiled value. The reason is that there is no longer an obvious choice of pseudo-expectation operator which will satisfy the requisite properties. In particular, the game value can no longer be described by a sum of quantities that look like $\langle\psi|A_{x}B_{y}|\psi\rangle$, as we no longer have the guarantee $[A_{x},B_{y}]=0$ afforded to us before by the spatial separation, meaning that the $g$ polynomial may not be zero when evaluated on the “valid” strategies for the compiled game. In [NZ23], a choice of pseudo-expectation operator was carefully defined so that the pseudo-expectation of $g(A_{x},B_{y})$ was zero by fiat, and (weaker versions of) properties (1) and (2) could be argued by reduction to cryptographic security properties. However, the fact that the measurement operators used by Alice in the compiled game return encrypted outcomes was an obstacle to carrying out these cryptographic reductions because the security only holds with respect to efficient distinguishers. In [NZ23], this obstacle was overcome in the CHSH game because the SOS certificate for this game has a particularly simple dependence on the $A_{x}$ observables. Building on a work of Slofstra [Slo11], we show that any XOR game has a “nice” SOS certificate (in which each $q_{i}$ has degree $1$ and depends on at most one $A_{x}$ variable) and that these “nice” SOS certificates yield bounds on the compiled value as well as the quantum value.
@@ 1.4 Related work
During the completion of this work, we became aware that the self-testing result for the Magic Square game was independently obtained separately in unpublished works by Thomas Vidick, and by Fermi Ma and Alex Lombardi, using similar techniques to ours.
@@ 2.1 Cryptography
We adopt several definitions from [NZ23].
@@
Here we follow [NZ23] in considering security against uniform adversaries, and indeed all of our reductions are uniform. We remark that we can also define security against non-uniform adversaries to obtain a stronger conclusion at the cost of relying on a stronger cryptographic assumption (specifically, QHE secure against non-uniform adversaries, which is quite standard in cryptography). We now recall the notion of quantum homomorphic encryption (QHE).
@@
    • Game 2: Start with a key $\mathrm{sk}\leftarrow\mathrm{Gen}(1^{\lambda})$, $c\in\mathrm{Enc}(\mathrm{sk},x)$, and $|\psi\rangle_{AB}$. Apply $\mathrm{Eval}$ with input $C$, register $A$, and ciphertext $c$ to obtain $\tilde{c}$. Compute $\tilde{y}\leftarrow\mathrm{Dec}(\mathrm{sk},\tilde{c})$. Return $\tilde{y}$ and register $B$.
@@ 2.2 Non-Local Games
A commuting operator strategy $\mathcal{S}$ for a nonlocal game $\mathcal{G}=(I_{A},I_{B},O_{A},O_{B},\pi,V)$ is a tuple $\mathcal{S}=(\mathcal{H},|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$, consisting of a Hilbert space $\mathcal{H}$, a state $|\psi\rangle\in\mathcal{H}$, and two collections of mutually commuting PVMs $\{A_{xa}\}_{a\in O_{A}}$ acting on $\mathcal{H}$ for each $x\in I_{A}$ for Alice and PVMs $\{B_{yb}\}_{b\in O_{B}}$ acting on $\mathcal{H}$ for each $y\in I_{B}$ for Bob, i.e. $[A_{xa},B_{yb}]=0$ for all $a,b,x,y\in O_{A}\times O_{B}\times I_{A}\times I_{B}$. Like for quantum strategies, we will often omit the Hilbert space and write $\mathcal{S}=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$ for a commuting operator strategy.
We can also define the commuting operator (also known as the quantum commuting) value of a nonlocal game $\omega_{qc}^{*}(\mathcal{G})=\sup_{S}\omega_{qc}(\mathcal{S},\mathcal{G})$ to be the supremum over all commuting operator strategies $S$ compatible with $\mathcal{G}$. It is not hard to see that every quantum (tensor) strategy is a commuting operator strategy. The converse holds if we restrict our commuting operator strategies to be finite dimensional (i.e. $\mathcal{H}$ is finite dimensional). However, there are examples of nonlocal games $\mathcal{G}$ for which there is a perfect (wins with probability $1$) commuting operators strategy but no perfect tensor-product strategy, see for example [Slo16].
@@ 2.3 XOR Games
Given a commuting operator strategy for an XOR game $\mathcal{S}=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$ we can derive an expression for the commuting operator bias of the strategy $\beta_{qc}(\mathcal{S},\mathcal{G})$ in terms of ±$\pm 1$-valued observables (i.e. self-adjoint unitaries) as follows:
$1\displaystyle 2\omega_{qc}(\mathcal{S},\mathcal{G})-$ 1 $=2\left[\sum_{x,y}\pi(x,y)\sum_{a,b}V(a,b,x,y)\langle\psi|A_{xa}B_{yb}|\psi\rangle\right]-1$ (2.1)
$=2\left[\sum_{x,y}\pi(x,y)\sum_{a,b}\frac{1}{2}(1+(-1)^{f(x,y)+a+b})\langle\psi|A_{xa}B_{yb}|\psi\rangle\right]-1$ (2.2)
$=\sum_{x,y}\pi(x,y)\sum_{a,b}(-1)^{f(x,y)+a+b}\langle\psi|A_{xa}B_{yb}|\psi\rangle$ (2.3)
$=\sum_{x,y}\pi(x,y)(-1)^{f(x,y)}\langle\psi|\left(\sum_{a}(-1)^{a}A_{xa}\right)\left(\sum_{b}(-1)^{b}B_{yb}\right)|\psi\rangle$ (2.4)
@@
where $A_{x}=A_{x0}-A_{x1}$ and $B_{y}=B_{y0}-B_{y1}$ are the ±$\pm 1$-valued (binary) observables for all $x,y\in I_{A}\times I_{B}$. Note that this construction is reversible, and thus strategies consisting of these ±$\pm 1$-valued observables are equivalent to commuting operator strategies for $\mathcal{G}$. With this in mind, we define the optimal commuting operator bias of $\mathcal{G}$ to be $\beta_{qc}^{*}(\mathcal{G})=2\omega_{qc}^{*}(\mathcal{G})-1$. Since all quantum strategies are commuting operator strategies we can similarly define the quantum bias $\beta_{q}(S,\mathcal{G})=2\omega_{q}(S,\mathcal{G})-1$. We define the optimal quantum bias of $\mathcal{G}$ to be $\beta_{q}^{*}(\mathcal{G})=2\omega_{q}^{*}(\mathcal{G})-1$.
@@
[…]
@@
The first two items are Theorem 3.1 and Corollary 3.2 of [Slo11], together with the fact that any optimal quantum strategy is an optimal vector strategy, which follows from Corollary 2.11. The formulation of the primal in (3) is a direct consequence of Corollary 2.11, and the duality of the semi-definite programs is shown in the proof of Theorem 3.1 of [Slo11]. ∎

### v3

@@ 1 Introduction
In a nonlocal game, a polynomially-bounded classical verifier interacts with multiple non-communicating provers. A set of predetermined rules governs whether the provers win or lose the game. The classical value of a nonlocal game is the highest winning probability that any classical (or unentangled) provers can achieve, whereas the quantum value is the highest winning probability that any quantum (potentially entangled) provers can achieve. Although the study of nonlocal games originated with Bell’s theorem [3], they have become central in the study of quantum complexity theory through their connection with multiprover interactive proofs [6]. This connection has led to many protocols for efficiently certifying classical and quantum computations in the two-prover setting [22, 8, 11, 13].
A central question in the nonlocal game paradigm is whether the techniques and results can be translated to a model where a verifier interacts with a single prover. The single-prover setting is theoretically appealing and practically motivated since a non-communication assumption between provers can be challenging to enforce. Recently, Kalai et al. [14] proposed a generic procedure that compiles any $k$-player one-round nonlocal game into a multi-round single-prover protocol, using a quantum homomorphic encryption scheme [16, 4]. In the same work, the authors showed that the compilation process preserves the classical value of the nonlocal game. Shortly after, Natarajan and Zhang [19] proved that the same compiler preserves the quantum value of the $2$-player CHSH game [5]. This was sufficient to achieve efficient delegation of BQP computations, however, their work left open the question of whether the Kalai et al. compiler preserved the quantum value for a more general class of games.
A natural generalization of the CHSH game is the class of two-player XOR games [6], where the winning predicate depends on the XOR of the players’ single-bit answers. These games have a foundational place in the theory of nonlocal games thanks to an influential theorem of Tsirelson [26, 27], which exactly characterizes their optimal quantum strategies in terms of a semidefinite program. This makes XOR games one of the few classes of nonlocal games where the optimal quantum strategies are well understood, and finding the quantum value is computationally tractable. As such, they are a convenient stepping stone towards understanding more general families of nonlocal games.
@@ 1.1 Results
Our main result is that the quantum value of any two-player XOR game is preserved by the compilation procedure of [14], resolving a question posed in [19]. We give an informal statement of our result; a more detailed statement can be found in Theorem 3.1.
@@
The negligible factors in Theorem 1.1 come from the fact that the compiled game assumes the existence of quantum homomorphic encryption, whose security is computational and depends on chosen parameters. It is well-known that quantum homomorphic encryption can be constructed from the hardness of the standard learning with errors (LWE) problem [15, 4], and thus our theorem holds up to the same assumption. A useful consequence of this result is that we can apply the extensive literature on XOR games to give new examples of well-behaved compiled games. For instance, we obtain the following new results as almost immediate corollaries of the proof of our main theorem:
  • Self-testing for all XOR games Using a result from Slofstra [24], we show that every compiled XOR game self-tests some properties of the provers’ measurement operators: in particular, for each XOR game there is a set of relations between the Alice and Bob operators that are approximately satisfied for any near-optimal strategy for the compiled game (Theorem 4.3). As a special case, we recover the self-test for anticommutation in [19].
  • Parallel repetition Building on [7], we show that the quantum value of compiled parallel repeated XOR games satisfies perfect exponential decay in the number of parallel repetitions (Theorem 5.5). This is the first parallel repetition theorem for compiled games.
@@
  • Self-testing for the Magic Square game: We show that the compilation of the Magic Square game of Mermin and Peres [18, 21, 1] is a self-test for a pair of anticommuting operators (Theorem 6.7). This mirrors the statement in the two-prover settings [28].
@@ 1.2 Technical Outline
Non-local games and the [14] compiler.
@@
It is easy to show that the compiled value is always at least the nonlocal value, as any nonlocal strategy can be implemented using the homomorphic encryption scheme. It is far less obvious how to establish upper bounds on the compiled value. This is because all known general techniques for upper-bounding the value of a nonlocal game make essential use of the spatial separation between the two provers. Nevertheless, we are able to show that the two specific techniques used by [19] to analyze the compiled CHSH game both generalize to all compiled XOR games. We believe the two approaches are complementary and improve our understanding of compiled nonlocal games from dual perspectives.
@@
An approach to upper-bounding the quantum value of a nonlocal game is to associate with a game a formal polynomial $p$, whose variables are identified with the family of observables $\{A_{x}\}$ and $\{B_{y}\}$ so that $\langle\psi|p\left(\{A_{x}\},\{B_{y}\}\right)|\psi\rangle$ is precisely the winning probability of the strategy. (For example, in the case of CHSH, the polynomial is $p\left(\{A_{x}\},\{B_{y}\}\right)=\frac{1}{2}+\frac{1}{8}(A_{0}B_{0}+A_{0}B_{1}+A_{1}B_{0}-A_{1}B_{1})$, and it can be easily verified that $\langle\psi|p\left(\{A_{x}\},\{B_{y}\}\right)|\psi\rangle$ captures the winning probability of the strategy $(\{A_{x}\},\{B_{y}\},|\psi\rangle)$ for the CHSH game. Moreover, to prove that the quantum value of a game is at most $\omega$, it suffices to give a sum-of-squares (SOS) certificate demonstrating that $\omega\mathds{1}-p\left(\{A_{x}\},\{B_{y}\}\right)$ is a positive operator whenever $p$ is evaluated on variables that corresponds to valid observables. Such a certificate consists of a decomposition of $\omega\mathds{1}-p\left(\{A_{x}\},\{B_{y}\}\right)$ as a positive operator
$\omega\mathds{1}-p\left(\{A_{x}\},\{B_{y}\}\right)=\sum_{i}q_{i}\left(\{A_{x}\},\{B_{y}\}\right)^{*}q_{i}\left(\{A_{x}\},\{B_{y}\}\right)+g\left(\{A_{x}\},\{B_{y}\}\right),$
@@
where the $q_{i}$’s and $g$ are polynomials, and $g$ is a polynomial that evaluates to zero whenever the variables are valid observables. In particular, $g$ lies in the ideal generated by a certain collection of polynomial constraints on the variables $A_{x}$ and $B_{y}$, including the commutator constraint $A_{x},B_{y}]=0$ for all $x,y$, which formally represents the spatial separation of the players.
@@
The reason why such an SOS certificate suffices to upper bound the game value in the nonlocal case is that (1) taking the expectation $\langle\psi|\omega\mathds{1}-p\left(\{A_{x}\},\{B_{y}\}\right)|\psi\rangle$ maps positive operators to positive numbers, and (2) $\langle\psi|p\left(\{A_{x}\},\{B_{y}\}\right)|\psi\rangle$ is equal to the winning probability of strategy $S=(\{A_{x}\},\{B_{y}\},|\psi\rangle)$. Property (1) gives that evaluating the expectation on $\omega\mathds{1}-p\left(\{A_{x}\},\{B_{y}\}\right)$ yields a positive number, and property (2) says that this positive number is equal to $\omega-p_{win}$, where $p_{win}$ is the winning probability of the strategy $S=(\{A_{x},B_{y}\},|\psi\rangle)$. Therefore, in the nonlocal case, taking the expectation plays the role of what is often known as a “pseudo-expectation” in the SOS literature.
@@
In general, an SOS certificate of the aforementioned form does not automatically yield an upper bound on the compiled value. The reason is that there is no longer an obvious choice of pseudo-expectation operator which will satisfy the requisite properties. In particular, the game value can no longer be described by a sum of quantities that look like $\langle\psi|A_{x}B_{y}|\psi\rangle$, as we no longer have the guarantee $A_{x},B_{y}]=0$ afforded to us before by the spatial separation, meaning that the $g$ polynomial may not be zero when evaluated on the “valid” strategies for the compiled game. In [19], a choice of pseudo-expectation operator was carefully defined so that the pseudo-expectation of $g\left(\{A_{x}\},\{B_{y}\}\right)$ was zero by fiat, and (weaker versions of) properties (1) and (2) could be argued by reduction to cryptographic security properties. However, the fact that the measurement operators used by Alice in the compiled game return encrypted outcomes was an obstacle to carrying out these cryptographic reductions because the security only holds with respect to efficient distinguishers. In [19], this obstacle was overcome in the CHSH game because the SOS certificate for this game has a particularly simple dependence on the $A_{x}$ observables. Building on a work of Slofstra [24], we show that any XOR game has a “nice” SOS certificate (in which each $q_{i}$ has degree $1$ and depends on at most one $A_{x}$ variable) and that these “nice” SOS certificates yield bounds on the compiled value as well as the quantum value. As a corollary, we obtain a one-sided rigidity or “self-testing” statement for any compiled XOR game: For any near-optimal strategy, it must hold that $A_{x}|\psi\rangle\approx\sum_{y}c_{xy}B_{y}A_{x}|\psi\rangle$ for explicit coefficients $c_{xy}$.
Although the approach using SOS allows us to obtain a bound on the compiled value for any XOR game, we mention that there is an alternative proof of this fact that employs Tsirelson’s correspondence in the compiled setting. Tsirelson’s correspondence shows that every vector strategy gives rise to a quantum strategy for the nonlocal XOR game [27, 7]. Here, by applying Jensen’s inequality to Bob’s operators, we can bound (up to a negligible function) the bias of every quantum strategy for the compiled XOR game by the bias achieved by some vector strategy for the nonlocal XOR game. Therefore, by Tsirelson’s correspondence we obtain a bound on the winning probability for the compiled game in terms of the optimal nonlocal value of the original XOR game. This approach was also used in  [19] for the CHSH game and inspired the generalization here. However, as this approach does not recover the proof of the “nice” SOS certificates for XOR games we defer the alternate proof to the Appendix A.
@@ 1.4 Related work
During the completion of this work, we became aware that the self-testing result for the Magic Square game in Section 6 was independently obtained in unpublished works by Thomas Vidick, and separately Fermi Ma and Alex Lombardi, using similar techniques to ours. Shortly after posting the first draft of this work, the independent work [2] appeared and obtained the same result for the value of compiled XOR games.
@@ 2.1 Cryptography
We adopt several definitions from [19].
@@
Here we follow [19] in considering security against uniform adversaries, and indeed all of our reductions are uniform. We remark that we can also define security against non-uniform adversaries to obtain a stronger conclusion at the cost of relying on a stronger cryptographic assumption (specifically, QHE secure against non-uniform adversaries, which is quite standard in cryptography). We now recall the notion of quantum homomorphic encryption (QHE).
@@
[…]
@@ 2.2 Non-Local Games
A commuting operator strategy $\mathcal{S}$ for a nonlocal game $\mathcal{G}=(I_{A},I_{B},O_{A},O_{B},\pi,V)$ is a tuple $\mathcal{S}=(\mathcal{H},|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$, consisting of a Hilbert space $\mathcal{H}$, a state $|\psi\rangle\in\mathcal{H}$, and two collections of mutually commuting PVMs $\{A_{xa}\}_{a\in O_{A}}$ acting on $\mathcal{H}$ for each $x\in I_{A}$ for Alice and PVMs $\{B_{yb}\}_{b\in O_{B}}$ acting on $\mathcal{H}$ for each $y\in I_{B}$ for Bob, i.e. $A_{xa},B_{yb}]=0$ for all $a,b,x,y\in O_{A}\times O_{B}\times I_{A}\times I_{B}$. Like for quantum strategies, we will often omit the Hilbert space and write $\mathcal{S}=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$ for a commuting operator strategy.
We can also define the commuting operator (also known as the quantum commuting) value of a nonlocal game $\omega_{qc}^{*}(\mathcal{G})=\sup_{S}\omega_{qc}(\mathcal{S},\mathcal{G})$ to be the supremum over all commuting operator strategies $S$ compatible with $\mathcal{G}$. It is not hard to see that every quantum (tensor) strategy is a commuting operator strategy. The converse holds if we restrict our commuting operator strategies to be finite dimensional (i.e. $\mathcal{H}$ is finite dimensional). However, there are examples of nonlocal games $\mathcal{G}$ for which there is a perfect (wins with probability $1$) commuting operators strategy but no perfect tensor-product strategy, see for example [25].
@@ 2.3 XOR Games
Given a commuting operator strategy for an XOR game $\mathcal{S}=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$ we can derive an expression for the commuting operator bias of the strategy $\beta_{qc}(\mathcal{S},\mathcal{G})$ in terms of $\pm 1$-valued observables (i.e. self-adjoint unitaries) as follows:
$2\omega_{qc}(\mathcal{S},\mathcal{G})-1$ $=2\left[\sum_{x,y}\pi(x,y)\sum_{a,b}V(a,b,x,y)\langle\psi|A_{xa}B_{yb}|\psi\rangle\right]-1$ (2.1)
$=2\left[\sum_{x,y}\pi(x,y)\sum_{a,b}\frac{1}{2}(1+(-1)^{g(x,y)+a+b})\langle\psi|A_{xa}B_{yb}|\psi\rangle\right]-1$ (2.2)
$=\sum_{x,y}\pi(x,y)\sum_{a,b}(-1)^{g(x,y)+a+b}\langle\psi|A_{xa}B_{yb}|\psi\rangle$ (2.3)
$=\sum_{x,y}\pi(x,y)(-1)^{g(x,y)}\langle\psi|\left(\sum_{a}(-1)^{a}A_{xa}\right)\left(\sum_{b}(-1)^{b}B_{yb}\right)|\psi\rangle$ (2.4)
@@
where $A_{x}=A_{x0}-A_{x1}$ and $B_{y}=B_{y0}-B_{y1}$ are the $\pm 1$-valued (binary) observables for all $x,y\in I_{A}\times I_{B}$. Note that this construction is reversible, and thus strategies consisting of these $\pm 1$-valued observables are equivalent to commuting operator strategies for $\mathcal{G}$. With this in mind, we define the optimal commuting operator bias of $\mathcal{G}$ to be $\beta_{qc}^{*}(\mathcal{G})=2\omega_{qc}^{*}(\mathcal{G})-1$. Since all quantum strategies are commuting operator strategies we can similarly define the quantum bias $\beta_{q}(S,\mathcal{G})=2\omega_{q}(S,\mathcal{G})-1$. We define the optimal quantum bias of $\mathcal{G}$ to be $\beta_{q}^{*}(\mathcal{G})=2\omega_{q}^{*}(\mathcal{G})-1$.
@@
    $\begin{aligned} (P):\quad\mathop{\mathrm{max}}_{\begin{subarray}{c}u_{x},v_{y}\in\mathbb{R}^{d}\\ \|u_{x}\|=\|v_{y}\|=1\end{subarray}}\sum_{x\in I_{A},y\in I_{B}}G_{xy}\langle u_{x}|v_{y}\rangle\end{aligned}\qquad\quad\begin{aligned} (D):\quad\min_{\{r_{x}\},\{c_{y}\}}&\frac{1}{2}\sum_{x}r_{x}+\frac{1}{2}\sum_{y}c_{y}:\\ &\begin{pmatrix}\Delta(\mathbf{r})&-G^{T}\\ -G&\Delta(\mathbf{c})\end{pmatrix}\succeq 0,\end{aligned}$
@@
The first two items are Theorem 3.1 and Corollary 3.2 of [24], together with the fact that any optimal quantum strategy is an optimal vector strategy, which follows from Corollary 2.11. The formulation of the primal in (3) is a direct consequence of Corollary 2.11, and the duality of the semi-definite programs is shown in the proof of Theorem 3.1 of [24]. ∎